OpenStudy (anonymous):

Which shows three ordered pairs that are all solutions of the equation y = x + 7? A. (0, 7), (2, 9), and (3, 11) B. (0, 7), (2, 9), and (3, 10) C. (1, 7), (2, 9), and (3, 10) D. (1, 8), (2, 9), and (3, 11)

OpenStudy (xapproachesinfinity):

plug in the ordered pairs and which satisfy the equation

OpenStudy (anonymous):

@xapproachesinfinity thanks but i know how to do that, i just cant find a graph up to 11 :(

OpenStudy (xapproachesinfinity):

why do you need the graph?

OpenStudy (xapproachesinfinity):

i said plug the number into the equation to verify with ordered pairs work not graphing

OpenStudy (xapproachesinfinity):

take the first one for example we need to check if those are solutions we check all the three of them (0,7) 7=0+7 correct (2,9) 9=2+7=9 correct (3,11) 11=3+7=10 incorrect so the first is not good

OpenStudy (anonymous):

ok

OpenStudy (xapproachesinfinity):

all of the three have to be correct for them to be the solution of that equation

OpenStudy (xapproachesinfinity):

you need to do the same with other options